# Jiakang Yuan1,2, Xiangchao Yan2, Botian Shi2, Tao Chen1,🖂, Wanli Ouyang2 Bo Zhang2,‡,🖂, Lei Bai2,🖂, Yu Qiao2, Bowen Zhou2

### v2

@@ Abstract
The scientific research paradigm is undergoing a profound transformation owing to the development of Artificial Intelligence (AI). Recent works demonstrate that various AI-assisted research methods can largely improve research efficiency by improving data analysis, accelerating computation, and fostering novel idea generation. To further move towards the ultimate goal (i.e., automatic scientific research), in this paper, we propose Dolphin, the first closed-loop open-ended auto-research framework to further build the entire process of human scientific research. Dolphin can generate research ideas, perform experiments, and get feedback from experimental results to generate higher-quality ideas. More specifically, Dolphin first generates novel ideas based on relevant papers which are ranked by the topic and task attributes. Then, the codes are automatically generated and debugged with the exception-traceback-guided local code structure. Finally, Dolphin automatically analyzes the results of each idea and feeds the results back to the next round of idea generation. Experiments are conducted on the benchmark datasets of different topics and results show that Dolphin can generate novel ideas continuously and complete the experiment in a loop. We highlight that Dolphin can automatically propose methods that are comparable to the state-of-the-art in some tasks such as 3D point classification. Our homepage: https://alpha-innovator.github.io/Dolphin-project-page
🖂 Corresponding Authors, $\ddagger$ Project Leader
@@ 1 Introduction
The fast evolution of Artificial Intelligence (AI) [1, 2, 46, 10] has brought about a profound transformation in various fields  [48, 16, 24]. In the landscape of scientific research, AI modes are developed and play an important role in accelerating key research processes, including scientific data collection and processing [11], scientific computation [17, 7], and scientific innovation [27]. Under this trend, the research paradigm is shifting from completely human-driven research to AI-assisted research [3]. More recently, the continuous iteration and upgrading of LLMs has promoted the evolution of AI-assisted scientific research to automatic scientific research [23, 34].
@@
Figure 1: Comparisons of the four steps in the evolutionary trajectory towards auto-research including (a) Entirely human-driven research, (b) AI-assisted research, (c) Semi-automatic research, and (d) Auto-research.
@@
The evolutionary trajectory from human-driven research to automatic research consists of four stages as shown in Fig. 1. 1) The entirely human-driven stage requires humans to complete all aspects including idea generation and experiments. 2) In the AI-assisted research stage, researchers use LLMs-based tools [3, 37] to improve the research efficiency. For example, GPT-researcher [3] can help us to decompose complex tasks and generate research reports using LLMs. 3) The semi-automatic research stage enables automation in certain aspects of scientific research. Recently, several works [34, 20, 36] use LLMs to generate ideas for different topics based on relevant works automatically. 4) The ultimate stage of AI development in science is the automatic scientific research stage, where AI system can automatically conduct the entire research process from conception to completion. Recently, AI-Scientist [23] proposes an open-ended auto-research framework capable of performing the complete scientific workflow, from idea generation to experimental validation and academic paper writing.
@@
Despite the encouraging progress made in existing works, several fundamental challenges continue to hinder the advancement of automatic scientific research. First, current studies face significant difficulties in accurately assessing the effectiveness of AI-generated ideas. Most existing works [34, 20, 36] either rely on human studies or directly employ LLMs to evaluate the quality (e.g., novelty) of generated ideas. However, merely focusing on the novelty of an idea itself does not adequately reflect its effectiveness in experimental validation. While works like AI-Scientist [23] consider experimental validation, they evaluate on the self-constructed simple datasets, making it challenging to draw meaningful comparisons with existing methods in the same field. Second, a critical limitation in previous works [34, 20, 36] is the absence of a feedback mechanism between the experimental validation unit and the idea generation unit – a process that is fundamental to human research. Human researchers typically iterate their ideas or refine existing approaches based on experimental outcomes, which serves as a crucial pathway for improving the quality of research ideas.
@@
To address these challenges and facilitate further progress towards automatic scientific research, in this work, we propose Dolphin, the first closed-loop open-ended automatic scientific research framework which is composed of three key steps in the research cycle (i.e., idea generation, experimental verification, and results feedback). In each research loop, Dolphin first retrieves the papers related to the given topic and generates novel and independent ideas. Then, the codes can be generated and debugged under the guidance of the experimental plan, which is directly generated by LLMs. Finally, Dolphin automatically analyzes the experimental results of successfully executed experiments which will guide the next loop of idea generation. Besides, for the two key steps in the auto-research loop (i.e., idea generation and experiment verification), we claim that the quality of ideas and execution rate of codes are crucial for enhancing the efficiency of auto-research. To improve the quality of ideas, inspired by ResearchAgent [4], we find that the relevance of retrieved papers to the topic is positively correlated with the quality of generated ideas. Therefore, instead of directly using the retrieved papers as references for idea generation, Dolphin filters these papers by judging the topic relevance and task attribute relevance between the retrieved papers and the input topic. Further, to improve the execution rate of Dolphin, we design an error-traceback-guided debugging process that analyzes the local code structure related to the error-traceback to guide the debugging process.
To further evaluate the effectiveness of automatic research, we conduct experiments on the benchmarks such as ModelNet40 [45], CIFAR-100 [19], and SST-2 [35], covering different tasks like 2D image classification, 3D point classification and sentiment classification. Our findings reveal that Dolphin generates ideas that boost the performance on these benchmarks compared to selected baselines like PointNet [28], WideResNet [12], and BERT-base [8]. We would like to emphasize that, Dolphin can propose methods based on PointNet that achieve performance comparable to human-designed state-of-the-art 3D classification methods. Besides, the quality of generated ideas improves through feedback, validating the effectiveness of the proposed closed-loop design.
@@
  • We propose Dolphin, the first closed-loop open-ended automatic research framework that covers three key steps in the human research cycle, including generating ideas, performing experiments, and feedback.
  • To improve the efficiency of automatic research, we propose task-attribute-guided paper ranking and exception-traceback-guided debugging process to improve the quality of generated ideas and the successful rate of code execution, respectively.
  • Experimental results on benchmark datasets show that Dolphin can generate high-quality ideas and perform experiments in the loop. We were supervised to observe that Dolphin is capable of achieving performance comparable to state-of-the-art methods in certain areas through automated research.
@@ 2 Related Works
2.1 AI for Automatic Scientific Research
@@
Open-ended Scientific Research. Recent studies [26, 42, 47, 27, 20, 23, 13] have demonstrated that Large Language Models (LLMs) have the potential to generate novel research ideas, a finding that has sparked widespread discussion in academia. Yang et al. [47] focuses on the social science domain and develops a framework called MOOSE based on LLM prompting, incorporating diverse feedback mechanisms to enhance the quality of generated ideas. Li et al. [20] introduces an innovative chain structure in a research domain, aiming to enhance the reasoning capabilities of LLMs. Besides, some iterative optimization strategies [42, 13] can also improve the novelty of ideas. While these approaches strive to improve the novelty of scientific ideas, a crucial aspect is the empirical validation of their practical effectiveness. AI-Scientist [23] proposes an end-to-end framework that automates the entire process from idea generation to experimental execution and paper writing. However, the experimental validation of its idea generation remains preliminary and lacks evaluation on real datasets or scenarios. Besides, the framework lacks a feedback mechanism from experimental validation to idea generation, unlike human researchers who iteratively refine their hypotheses based on experimental outcomes.
@@
Scope-limited Scientific Research. Several studies have successfully applied LLMs to specific scientific discovery tasks. For example, AutoML-GPT [52] leverages LLMs for hyperparameter tuning by combining model and data descriptions as prompts to predict training logs. AgentHPO [21] proposes a creator-executor framework that conducts experiments for specific hyperparameters and iteratively optimizes them based on historical trials. Similarly, MLCopilot [51] constructs an experience pool from historical data to enable LLMs-based hyperparameter prediction. EvoPrompting [6] improves the in-context prompting examples of LLMs to achieve effective code-level neural architecture design. In contrast to these scope-limited approaches, our method focuses on more open-ended, autonomous scientific discovery, encompassing the entire process from idea generation to final experimental validation, thereby achieving a truly closed-loop logic across the complete research lifecycle.
2.2 LLMs-assisted Research Tools
As LLMs [1, 46, 5, 10, 2, 38] continue to evolve rapidly, they have emerged as powerful tools for boosting research efficiency. They have been utilized such as report or survey generation [3, 44, 14], code development [37, 15, 53, 50], and data analysis [32, 18]. A notable example is GPT-Researcher [3], an autonomous agent based on LLMs and multi-agent systems to generate comprehensive research reports on the specified topics by effectively leveraging external knowledge bases. In the code development domain, tools such as GitHub Copilot [37] leverage the power of LLMs to provide intelligent, context-aware code suggestions and completions within Integrated Development Environments (IDEs), which significantly streamline the development process and enhance programming efficiency. The integration of LLMs as research assistants [3] opens up promising avenues for automating routine tasks, accelerating the overall research workflow. This technological breakthrough represents a significant step forward in augmenting human research capabilities.
@@ 3 Methods
Figure 2: The overview of Dolphin, an closed-loop open-ended auto-research framework. Dolphin first generates a set of ideas based on the retrieved papers. After filtering ideas, experimental plans will be generated for these filtered ideas. Then, codes can be generated and debugged using the proposed error-traceback-guided debugging process. Finally, the results of successfully executed experiments will be auto-analyzed and reflected into the next round of ideas generation.
In this section, we introduce Dolphin, a closed-loop open-ended auto-research framework as shown in Fig. 2, which is mainly composed of an idea generation process, an experiments verification process, and a result feedback process. The closed-loop means that the experimental results will be fed back into the idea generation process and the above three processes form a research cycle. In the idea generation process, Dolphin first retrieves papers based on the input topic and then filters papers that are not relevant to the topic. Then, the retrieved papers serve as references to guide LLMs to generate ideas, which will be further filtered through novelty and independence check process. Subsequently, Dolphin formulates experimental plans and proceeds to generate and debug code using a specifically designed error-traceback-guided debugging process. Finally, the results will be analyzed and utilized as feedback for the next cycle of ideas generation. In the following section, we will detail the ideas generation process, experiments verification process, and results feedback process in Sec. 3.1, Sec. 3.2, and Sec. 3.3 respectively.
@@ 3.1 Ideas Generation Process
“A good beginning is half done.” As the beginning of the research cycle, high-quality ideas are crucial to the entire research. A promising approach to generating high-quality ideas is to imitate the behaviors of human researchers. They typically first conduct literature reviews and then generate ideas based on the literature [31]. To generate novel ideas, Dolphin typically divided the idea generation process into two steps including 1) paper retrieval and ranking, and 2) ideas generation and filtering.
@@
Paper Retrieval and Ranking. To generate high-quality ideas, the first step is to retrieve papers that are relevant to the topic. Given a research topic (e.g., 3D classification), Dolphin begins by searching for relevant papers using Semantic Scholar API [^1] https://www.semanticscholar.org/product/api, obtaining the essential information such as titles and abstracts. However, the initially retrieved papers are not always directly related to the input topic, which can limit their usefulness in generating subsequent ideas. For instance, if the input topic is 3D classification, some retrieved papers might pertain to 3D detection [44]. Although these topics are interconnected, they typically focus on different challenges. As a result, it is necessary to filter out papers that are irrelevant to the specific topic.
@@
To this end, we design a paper ranking process that aims to assign a higher score to the paper relevant to the input topic. In detail, Dolphin ranks the retrieved papers based on two main criteria: 1) relevance to the input topic, and 2) alignment of task attributes with those of the input topic. The task attributes typically define a task, including model inputs, model outputs, and other characteristics. Specifically, the LLM is first utilized to extract the task attributes of the input topic and then prompt it to score (i.e., 1-10) each retrieved paper based on the designed criteria. The detailed prompts can be found in our supplementary material and we use gpt-4o-2024-08-06 to retrieve and score papers at this stage. After scoring, Dolphin filters out papers with scores below 8, using the remaining papers as the references of the subsequent ideas generation process.
@@
Ideas Generation and Filtering. After obtaining retrieved papers that are related to the topic, the next step is to generate novel ideas based on the references. Dolphin begins by prompting the LLM to generate $N$ novel and non-redundant ideas, each comprising a title, a brief experiment plan, and a summary. However, these generated ideas are not consistently novel, and some ideas are similar to one another. As a result, directly performing experiments on such ideas will cost substantial time and computational resources, further reducing research efficiency. To address this, we introduce a further idea-filtering procedure that filters out non-novel or redundant ideas.
@@
To be specific, Dolphin first examines the independence of ideas to ensure non-redundancy. Given $N$ ideas $[I_{1},I_{2},...,I_{N}]$, Dolphin first extract the sentence-level embedding $[f_{1},f_{2},...,f_{N}]$ based on the summary of each ideas. Then, an idea bank $\mathbf{B}$ is constructed to store embeddings of the ideas that have been checked to be independent. $\mathbf{B}$ is initialized as empty in the first loop and initialized with previous ineffective ideas in the following loops which will be introduced in Sec. 3.3. When examining the $i$-th idea, Dolphin calculates its cosine similarity with existing ideas stored in $\mathbf{B}$ and discards the current idea when the max similarity exceeds a predetermined threshold (i.e., 0.8). After filtering out redundant ideas, the remaining $N^{\prime}$ ideas will undergo the novelty check. Following AI-Scientist [23], Dolphin simply prompt the LLMs to decide whether the idea is novel based on the searched papers by Semantic Scholar API1. Only ideas identified as novel will proceed to the subsequent experimental verification process.

### v3

# Dolphin: Moving Towards Closed-loop Auto-research through Thinking, Practice, and Feedback

@@ Abstract
The scientific research paradigm is undergoing a profound transformation owing to the development of Artificial Intelligence (AI). Recent works demonstrate that various AI-assisted research methods can largely improve research efficiency by improving data analysis, accelerating computation, and fostering novel idea generation. To further move towards the ultimate goal (i.e., automatic scientific research), in this paper, we introduce Dolphin, a closed-loop LLM-driven framework to enhance the automation level of scientific research. Dolphin first generates novel ideas based on feedback from previous experiments and relevant papers ranked by the topic and task attributes. Then, the generated ideas can be implemented using a code template refined and debugged with the designed exception-traceback-guided local code structure. Finally, Dolphin automatically analyzes the results of each idea and feeds the results back to the next round of idea generation. Experiments are conducted on the benchmark datasets of different topics and a subset of MLE-bench. Results show that Dolphin can continuously improve the performance of the input topic in a loop. We highlight that Dolphin can automatically propose methods that are comparable to the state-of-the-art in some tasks such as 3D point classification.
Dolphin: Moving Towards Closed-loop Auto-research through Thinking, Practice, and Feedback
Jiakang Yuan♣,♠, Xiangchao Yan, Shiyang Feng♣,♠, Bo Zhang♠,‡,🖂, Tao Chen♣,🖂, Botian Shi, Wanli Ouyang, Yu Qiao, Lei Bai♠,🖂, Bowen Zhou Fudan University, Shanghai Artificial Intelligence Laboratory [Uncaptioned image] https://alpha-innovator.github.io/Dolphin-project-page/ [Uncaptioned image] https://github.com/Alpha-Innovator/Dolphin
🖂 Corresponding Authors, $\ddagger$ Project Lead
@@ 1 Introduction
The rapid evolution of AI Achiam et al. (2023); Anthropic (2024); Yang et al. (2024a); Dubey et al. (2024) has profoundly transformed various fields  Yao et al. (2024); Yan et al. (2025); Luo et al. (2022), accelerating scientific research processes like scientific data collection and processing Han et al. (2024); Yan et al. (2025), computation Jumper et al. (2021); Chen et al. (2023), and innovation Qi et al. (2024). Under this trend, the research paradigm is shifting from completely human-driven research to AI-assisted research Assafelovic (2023). More recently, the continuous upgrading of LLMs has promoted the evolution of AI-assisted research to automatic scientific research Lu et al. (2024); Si et al. (2024).
@@
Figure 1: Comparisons of the four stages in the evolutionary trajectory towards auto-research including (a) Entirely human-driven research, (b) AI-assisted research, (c) Semi-automatic research, and (d) Auto-research.
@@
The evolutionary trajectory from human-driven research to automatic research consists of four stages as shown in Fig. 1. 1) The entirely human-driven stage requires humans to complete all aspects including idea generation and experiments. 2) In the AI-assisted research stage, researchers use LLMs-based tools Assafelovic (2023); Team (2023a) to improve the research efficiency. For example, GPT-researcher Assafelovic (2023) can help us to decompose complex tasks and generate research reports using LLMs. 3) The semi-automatic research stage enables automation in certain processes of scientific research. For example, recent works Si et al. (2024); Li et al. (2024); Su et al. (2024) leverage LLMs to automatically generate ideas for different topics by drawing on relevant works. 4) The fourth stage of AI-promoting scientific research is auto-research stage, where AI automatically handles the entire research process from conception to completion. Recently, AI-Scientist Lu et al. (2024) introduced an open automatic research framework automating key tasks including idea generation, experimental verification, and academic paper writing.
Despite the encouraging progress made in existing works, auto-research still faces key challenges:
@@
First, accurately assessing the effectiveness of AI-generated ideas remains a major hurdle. Most studies Si et al. (2024); Li et al. (2024); Su et al. (2024) rely on either human evaluation or LLMs to evaluate the quality of generated ideas. However, merely focusing on the novelty of an idea itself does not adequately reflect its effectiveness in experimental validation. While works like AI-Scientist Lu et al. (2024) incorporate experimental validation, they use simple, the self-constructed datasets, limiting meaningful comparisons with existing methods in the same field.
Second, another limitation of prior works Si et al. (2024); Li et al. (2024); Su et al. (2024) lies in the absence of a feedback mechanism between the experimental validation and the idea generation – a process that is fundamental to human research. Human researchers refine their ideas and approaches iteratively based on experimental outcomes, which serves as a crucial pathway for improving the quality of research ideas.
@@
To address these challenges and facilitate further progress towards automatic scientific research, in this work, we propose Dolphin, a closed-loop auto-research framework which is composed of three key stages in the research cycle (i.e., idea generation, experimental verification, and results feedback). In each research loop, Dolphin first generates ideas based on the previous experimental results and the retrieved relevant papers which are filtered by judging the topic relevance and task attribute relevance between the retrieved papers and the input topic. Then, the proposed idea can be further implemented using a code template refined and debugged by the designed traceback-guided debugging module, which analyzes the local code structure related to the error-traceback to streamline the debugging process. Finally, Dolphin automatically analyzes the outcomes of successfully executed experiments, providing guidance for the next iteration of idea generation.
To experimentally validate the generated scientific ideas, we selected commonly-used public benchmarks across different tasks and modalities such as ModelNet40 Wu et al. (2015), CIFAR-100 Krizhevsky et al. (2009), and SST-2 Socher et al. (2013). The results demonstrate that Dolphin generates ideas that outperform the selected baselines like PointNet Qi et al. (2017a), WideResNet He et al. (2016), and BERT-base Devlin et al. (2019). We were supervised to observe that Dolphin is capable of proposing more concise method implementations while surpassing the performance of current human-designed state-of-the-art approaches in certain fields, as reported in Tab. 8. Besides, experiments on tasks from MLE-bench reveal that Dolphin can integrate with code generation pipelines like AIDE and further support version updates for technologies or code.
@@
  • We propose Dolphin, a closed-loop auto-research framework covering three key stages in the research cycle, including generating ideas, performing experiments, and feedback.
  • To improve the efficiency of auto-research, we propose task-attribute-guided paper ranking and exception-traceback-guided debugging process to improve the quality of generated ideas and the successful rate of code execution, respectively.
  • Experimental results on benchmark datasets show that Dolphin can generate high-quality ideas and conduct validations in the loop.
@@ 2 Related Works
Figure 2: Dolphin first generates a set of ideas based on the retrieved papers. After filtering ideas, experimental plans will be generated for these filtered ideas. Then, codes can be generated and debugged using the proposed error-traceback-guided debugging process. Finally, the results of successfully executed experiments will be auto-analyzed and reflected into the next round of ideas generation.
2.1 Open-ended Scientific Research
@@
Recent studies (Qi et al., 2023; Wang et al., 2024b; Yang et al., 2024b; Qi et al., 2024; Lu et al., 2024; Hu et al., 2024) have demonstrated that Large Language Models (LLMs) have the potential to generate novel research ideas, a finding that has sparked widespread discussion in academia. Li et al. (2024) generate ideas based on the analysis of the paper chain. Si et al. (2024) find that LLM-generated ideas are innovative but their feasibility needs improvement. Some works (Wang et al., 2024b; Hu et al., 2024) improve the novelty of ideas using iterative optimization strategies. In addition, several works (Qiu et al., 2023; Zhou et al., 2024) try to use LLMs to generate hypotheses. For example, Yang et al. (2023) generate hypotheses from web corpus and  Wang et al. (2023) generate hypotheses from literature. However, these works lack empirical verification of their practical effectiveness. Recently, AI-Scientist (Lu et al., 2024) and Agent Laboratory (Schmidgall et al., 2025) introduce the end-to-end framework that automates the process from idea generation to experimental execution and paper writing. However, the experimental validation of its idea generation remains limited, lacking evaluation conducted on real-world datasets or scenarios. Furthermore, the framework lacks a feedback mechanism that connects experimental validation to idea generation, unlike human researchers who iteratively refine their hypotheses based on experimental outcomes.
2.2 Scope-limited Scientific Research.
@@
Several studies have effectively applied LLMs to specific scientific discovery tasks. For example, AutoML-GPT (Zhang et al., 2023b) leverages LLMs for hyperparameter tuning by combining model and data descriptions as prompts to predict training logs. AgentHPO (Liu et al., 2024) introduces a creator-executor framework to iteratively optimize hyperparameters based on historical trials. Similarly, MLCopilot (Zhang et al., 2023a) constructs an experience pool from historical data to enable LLMs-based hyperparameter prediction. EvoPrompting (Chen et al., 2024) improves the in-context prompting examples of LLMs to achieve effective code-level neural architecture design. In contrast to these scope-limited approaches, our method focuses on more open-ended scientific discovery, spanning from idea generation to experimental validation, and achieving a fully closed-loop research lifecycle.
@@ 3 Methods
In this section, we introduce Dolphin, a closed-loop auto-research framework as shown in Fig. 2, which is mainly composed of an idea generation process, an experiments verification process, and a result feedback process. The closed-loop means that the experimental results will be fed back into the idea generation process and the above three processes form a research cycle. In the idea generation process, Dolphin generates ideas based on the feedback and retrieved papers that are filtered by a designed task-attribute-guided paper ranking process. Then, Dolphin formulates experimental plans and proceeds to generate and debug code using a specifically designed error-traceback-guided debugging process. Finally, the results will be analyzed and utilized as feedback for the next cycle of ideas generation. The following sections detail the ideas generation process, experiments verification process, and results feedback process in Sec. 3.1, Sec. 3.2, and Sec. 3.3 respectively.
@@ 3.1 Ideas Generation Process
“A good beginning is half done.” As the beginning of the research cycle, high-quality ideas are crucial to the entire research. A promising approach to generating high-quality ideas is to imitate the behaviors of human researchers. They typically first conduct literature reviews and then generate ideas based on the literature Randolph (2019) and previous experimental experience (detail in Sec. 3.3). Dolphin typically divided the idea generation process into two steps including 1) paper retrieval and ranking, and 2) ideas generation and filtering.
@@
Paper Retrieval and Ranking. To generate high-quality ideas, the first step is to retrieve papers that are relevant to the topic. Given a research topic (e.g., 3D classification), Dolphin begins by searching for relevant papers using Semantic Scholar API [^1] https://www.semanticscholar.org/product/api, obtaining the essential information such as titles and abstracts. However, the initially retrieved papers are not always directly related to the input topic, which can limit their usefulness in generating subsequent ideas. For instance, if the input topic is 3D classification, some retrieved papers might pertain to 3D detection Wang et al. (2024c). Although these topics are interconnected, they typically focus on different challenges. As a result, it is necessary to filter out papers that are irrelevant to the specific topic.
@@
To this end, we design a task-attribute-guided paper ranking process that aims to assign a higher score to the paper relevant to the input topic and task attribute. In detail, Dolphin ranks the retrieved papers based on two main criteria: 1) relevance to the input topic, and 2) alignment of task attributes with those of the input topic. The task attributes typically define a task, including model inputs, model outputs, and other characteristics. Specifically, the LLM is first utilized to extract the task attributes of the input topic and then prompt it to score (i.e., 1-10) each retrieved paper based on the designed criteria. The detailed prompts can be found in our appendix. After scoring, Dolphin filters out papers with scores below 8, using the remaining papers as the references of the subsequent ideas generation process.
@@
Ideas Generation and Filtering. After obtaining retrieved papers, the next step is to generate ideas based on the references. Dolphin begins by prompting the LLM to generate $N$ novel and non-redundant ideas, each comprising a title, a brief experiment plan, and a summary. However, these generated ideas are not consistently novel, and some ideas are similar to one another. As a result, directly performing experiments on such ideas will cost substantial time and computational resources, further reducing research efficiency. To address this, we introduce a further idea-filtering procedure that filters out non-novel or redundant ideas.
@@
To be specific, Dolphin first examines the independence of ideas to ensure non-redundancy. Given $N$ ideas $[I_{1},I_{2},...,I_{N}]$, Dolphin first extract the sentence-level embedding $[f_{1},f_{2},...,f_{N}]$ based on the summary of each ideas. Then, an idea bank $\mathbf{B}$ is constructed to store embeddings of the ideas that have been checked to be independent. $\mathbf{B}$ is initialized as empty in the first loop and initialized with previous ineffective ideas in the following loops which will be introduced in Sec. 3.3. When examining the $i$-th idea, Dolphin calculates its cosine similarity with existing ideas stored in $\mathbf{B}$ and discards the current idea when the max similarity exceeds a predetermined threshold (i.e., 0.8). After filtering out redundant ideas, the remaining $N^{\prime}$ ideas will undergo the novelty check. Following AI-Scientist Lu et al. (2024), Dolphin simply prompt the LLMs to decide whether the idea is novel based on the searched papers by Semantic Scholar API1. Only ideas identified as novel and independent will proceed to the subsequent experimental verification process.